# AI-Accelerated Flow Simulation: A Robust Auto-Regressive Framework for Long-Term CFD Forecasting

### v3

@@ 4.4 Application of proposed adaptive multi-step rollout
Performance of the three adaptive weighting strategies across different time integration schemes is assessed using a focused evaluation approach. To better evaluate the models’ ability to capture complex wake oscillatory vortex behavior, MSE results from this section are calculated based on $x$-velocity data from seven strategic probe points in the wake region (Fig. 9), rather than across the entire field. This targeted analysis provides precise assessment of how well each method captures critical vortex shedding dynamics.
@@
Results are summarized in Table 5. For direct prediction, the vanilla approach with fixed weights (MSE: 0.011) outperforms adaptive variants, likely because the direct prediction’s inherent simplicity does not benefit from complex weighting schemes. However, for derivative-based time integration methods, adaptive approaches show clear advantages. Both AW2 and AW3 demonstrate strong performance with forward Euler (MSE: 0.007 and 0.010 respectively), indicating that adaptive weighting helps balance short-term and long-term prediction accuracy more effectively than fixed weights. The most remarkable performance is achieved by combining Adams-Euler with AW3 (MSE: 0.002), which leverages both the superior temporal integration of Adams-Euler and the focused weighting strategy of AW3 that emphasizes the most critical prediction steps.
Table 5: MSE comparison for different adaptive weighting approaches across time integration schemes with $M=4$ multi-step rollout. MSE values calculated from seven probe points in Fig. 9.
@@
To understand how adaptive weights evolve during training, Fig. 10 shows weight evolution for AW2 with Adams-Euler. Two distinct phases emerge: (1) Stabilization phase (before 1000 epochs): minimal weight differences as the model learns basic temporal patterns; (2) Adaptation phase (after 1000 epochs): increased weight on the 4th loss term, indicating greater difficulty in long-term predictions. This trend validates the AW3 approach—since the first three loss terms show little distinction, emphasizing only the first and last steps captures the essential trade-off between immediate and long-term accuracy: evolution of adaptive weights under AW3 approach can be found in Fig. 15, B.
@@
To quantitatively evaluate the impact of AW3 improvements, we analyze temporal evolution of $x$-velocity at the seven probe points, comparing the two best-performing models (footnoted in Table 5): Model A (Forward Euler with AW2, MSE: 0.007) and Model B (Adams-Euler with AW3, MSE: 0.002).
Fig. 11 shows time-series predictions over the entire 350-snapshot forecasting horizon targeted in this study. Model A (Fig. 11(a)) generally captures vortex shedding patterns but shows noticeable discrepancies, especially at the 4th probe point with red colors, where amplitude and phase reproduction fails. Model B (Fig. 11(b)) closely matches ground truth across all probe points, successfully maintaining correct shedding period and amplitude throughout the entire prediction horizon. This demonstrates that Model B’s training strategy, Adams-Euler with AW3, enhances the model’s ability to capture both short-term and long-term temporal dynamics more effectively than treating all loss terms simultaneously.
@@
Figure 11: $x$-velocity time series at seven probe points over 350 future snapshots. From bottom to top: probe points 1-7 (Fig. 9). Solid lines: ground truth; circles: model predictions.
@@ 4.5 Comparison with conventional noise-injection approach for long-term rollout
We evaluate three model configurations (direct prediction, forward Euler, and Adams-Euler time integration schemes) trained with Gaussian noise $\mathcal{N}(0,0.16^{2})$ injected into input data. This noise level, identified as optimal in our previous work [27], is consistently applied during training without multi-step rollouts. We assess effectiveness using temporal behavior of $x$-velocity at seven probe points downstream of the cylinder. Figs. 12(a) and 12(b) present results for noise injection with forward Euler and Adams-Euler schemes respectively. For forward Euler, comparing adaptive multi-step rollout with AW2 (Fig. 11(a)) versus noise injection shows that noise injection exhibits inferior accuracy at all probe points, with oscillations deviating significantly from ground truth in amplitude and phase. This indicates noise injection alone is insufficient for maintaining long-term prediction accuracy. Similarly, for Adams-Euler, comparing AW3-based adaptive multi-step rollout (Fig. 11(b)) versus noise injection reveals less accurate reproduction of vortex shedding patterns, particularly for lower probe points and after 150 rollout steps.
@@
Figure 12: $x$-velocity time series at probe points over 350 future snapshots using noise injection. Solid lines: ground truth; circles: model predictions.
Comparing integration schemes with noise injection, Adams-Euler (Fig. 12(b)) still significantly outperforms forward Euler (Fig. 12(a)), further demonstrating the robust performance of the Adams-Bashforth time integration scheme across various training approaches.
Table 6 quantifies performance differences with MSE calculations for seven probe points. It’s important to note that noise injection itself provides substantial improvements over baseline approaches—comparing to Table 3 where no multi-step rollout or noise injection was applied (Direct: 0.125, Forward Euler: 0.138, Adams-Euler: 0.139), noise injection achieves significant performance gains (Direct: 0.019, Forward Euler: 0.017, Adams-Euler: 0.012), demonstrating its established effectiveness in the AR prediction community. However, for all time integration schemes, models trained with adaptive multi-step rollout consistently outperform even these improved noise injection results. The Adams-Euler scheme shows particularly dramatic enhancement from 0.012 with noise injection to 0.002 with adaptive multi-step rollout—representing not only substantial improvement over baseline methods but also a significant advance beyond the already powerful noise injection technique.
Table 6: MSE comparison between adaptive multi-step rollout, noise injection, and a combined approach. Each multi-step rollout configuration uses the best adaptive weighting approach (footnoted). MSE values calculated from seven probe points in Fig. 9.
@@
The performance gap stems from fundamental differences in addressing error accumulation. Noise injection builds resilience through input perturbations, making models more robust to small deviations, but does not explicitly address temporal dependencies and error propagation mechanisms inherent in AR predictions. In contrast, adaptive multi-step rollout enables learning from multiple future steps simultaneously, using strategically adjusted loss weights to focus on both immediate and distant predictions. This fundamental difference explains why adaptive multi-step rollout more accurately captures vortex shedding patterns compared to noise injection, as demonstrated in the time series comparisons (Figs. 11 and 12).
We also investigated combining multi-step rollout with noise injection techniques, which revealed contrasting behaviors as shown in Table 6. The combined approach yields the best performance for direct prediction (MSE: 0.008), while for derivative-based methods, this combination significantly degrades performance. This contrasting behavior can be attributed to how noise injection interacts with the model’s specific prediction target. For direct prediction, the model output is the exact snapshot for the next time step. Therefore, training with noise injection is highly effective, as it directly regularizes the model against the sort of input perturbations it will encounter when recursively using its own imperfect predictions during rollout. Conversely, for derivative-based methods such as Forward Euler and Adams-Euler, the model only predicts derivative values. The final future snapshot is then calculated through a time-integration formula that combines the model’s derivative output with the previous state (e.g., $\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\Delta t\cdot\frac{\delta\mathbf{u}}{\delta t}$). Because the model’s direct output is not the final prediction, the effect of noise injection on the input data is less direct. The integration step itself may act as a stabilizing buffer, making the additional regularization from noise injection less effective or even disruptive when combined with multi-step training.
4.6 Robustness evaluation under partial domain training
To further validate our framework’s robustness, we evaluate performance under partial domain training, where models are trained on only a spatial subset ($0.750.3<x<$ 0.75 and $0.3280.128<y<$ 0.328) of the original vortex shedding domain. This challenging scenario simulates practical applications with limited spatial coverage or memory constraints.
We test the four models that previously demonstrated satisfactory results from Table 5: direct prediction with vanilla multi-step rollout, forward Euler with AW2, forward Euler with AW3, and Adams-Euler with AW3. Figure 13 shows flow field predictions after 300 rollout steps: partial domain mesh used for training is visualized in Fig. 13(b)$\sim$13(e). Direct prediction fails completely, while forward Euler methods show limited success. Adams-Euler with AW3 demonstrates the best performance, maintaining high accuracy and successfully reproducing vortex shedding patterns within the constrained region.
@@ 4.6 Robustness evaluation under partial domain training
Quantitative analysis using seven probe points (Fig. 14) confirms these results. Direct prediction completely fails to capture vortex shedding oscillations (MSE: 0.019), while forward Euler methods show gradual degradation (MSE: 0.011-0.013). Adams-Euler with AW3 achieves the lowest error (MSE: 0.008), accurately maintaining vortex shedding frequency, amplitude, and phase throughout the entire 350-step prediction horizon. These results validate the exceptional robustness of our Adams-Bashforth time integration with adaptive multi-step rollout framework, confirming its broad applicability for practical engineering applications where complete spatial information may be unavailable.
@@ 5 Conclusions and future work
This study presents a comprehensive framework for enhancing long-term auto-regressive predictions in SciML models through the novel application of numerical time-integration schemes and adaptive multi-step rollout techniques. Our systematic evaluation across canonical 2D PDEs (advection, heat, and Burgers’ equations) first established a key hypothesis: as physical complexity increases, more sophisticated rollout techniques become essential for optimal performance. This trend was decisively validated in our most challenging test case of complex Navier-Stokes dynamics. For this system, our most advanced adaptive weighting strategies (AW2/AW3) proved crucial for achieving robust, long-term accuracy, confirming the insight gained from the simpler systems. By combining the two-step Adams-Bashforth scheme with these adaptive strategies, our lightweight GNN model—containing only 1,177 trainable parameters—demonstrated meaningful effectiveness under harsh constraints, achieving accurate predictions of complex Navier-Stokes dynamics across 350 future time steps and reducing the mean squared error from 0.125 to 0.002. Overall, our integrated methodology delivers an 89% improvement over fixed-weight approaches and outperforms standard noise injection by 83%, while maintaining robustness even on truncated meshes. This powerful yet resource-efficient framework is designed to be model-agnostic, ensuring these advancements can benefit diverse scientific domains without specialized adaptations.
Several promising directions can emerge for future research. First, investigating higher-order time integration schemes beyond the two-step Adams-Bashforth could further enhance prediction stability, particularly for systems with complex temporal dynamics. Second, developing more computationally efficient multi-step rollout strategies would reduce the gradient computation overhead inherent in training across multiple future steps. Third, incorporating domain-specific physical principles—such as conservation laws from the Navier-Stokes equations—directly into the framework could enhance both accuracy and physical consistency while potentially reducing data requirements [34, 35]. Finally, systematic evaluation across diverse scientific domains, from climate modeling to materials science, would provide deeper insights into the framework’s universal applicability and reveal domain-specific optimization opportunities, advancing toward more versatile and scalable AR prediction solutions.

### v4

@@ 4.4 Application of proposed adaptive multi-step rollout
Performance of the three adaptive weighting strategies across different time integration schemes is assessed using a focused evaluation approach. To better evaluate the models’ ability to capture complex wake oscillatory vortex behavior, MSE results from this section are calculated based on $x$-velocity data from seven strategic probe points in the wake region (Figure 9), rather than across the entire field. This targeted analysis provides precise assessment of how well each method captures critical vortex shedding dynamics.
@@
Results are summarized in Table 5. For direct prediction, the vanilla approach with fixed weights (MSE: 0.011) outperforms adaptive variants, likely because the direct prediction’s inherent simplicity does not benefit from complex weighting schemes. However, for derivative-based time integration methods, adaptive approaches show clear advantages. Both AW2 and AW3 demonstrate strong performance with forward Euler (MSE: 0.007 and 0.010 respectively), indicating that adaptive weighting helps balance short-term and long-term prediction accuracy more effectively than fixed weights. The most remarkable performance is achieved by combining Adams-Euler with AW3 (MSE: 0.002), which leverages both the superior temporal integration of Adams-Euler and the focused weighting strategy of AW3. This represents a significant 89% reduction in mean squared error compared to the conventional fixed-weight multi-step rollout (which had an MSE of 0.018), highlighting the practical benefit of our adaptive approach.
Table 5: MSE comparison for different adaptive weighting approaches across time integration schemes with $M=4$ multi-step rollout. MSE values calculated from seven probe points in Figure 9. Strouhal numbers (StSt) are also provided under each MSE value, where ground truth StSt value is 0.1438.
@@
To understand how adaptive weights evolve during training, Figure 10 shows weight evolution for AW2 with Adams-Euler. Two distinct phases emerge: (1) Stabilization phase (before 1000 epochs): minimal weight differences as the model learns basic temporal patterns; (2) Adaptation phase (after 1000 epochs): increased weight on the 4th loss term, indicating greater difficulty in long-term predictions. This trend validates the AW3 approach—since the first three loss terms show little distinction, emphasizing only the first and last steps captures the essential trade-off between immediate and long-term accuracy: evolution of adaptive weights under AW3 approach can be found in Figure 16, B.
@@
To quantitatively evaluate these improvements, we analyze the temporal evolution of the predictions by Model A (forward Euler with AW2) and B (Adams-Euler with AW3) selected in Table 5. Figure 11 shows the 350-snapshot forecast, with the corresponding time-varying MSE plots providing a measure of error accumulation. Model A’s predictions (Figure 11(a)) show noticeable discrepancies in amplitude and phase. This is confirmed by its time-varying MSE plot (Figure 11(c)), which reveals a clear upward trend indicating steady error accumulation that reaches a maximum MSE of approximately 0.0125. In stark contrast, Model B’s predictions (Figure 11(b)) closely match the ground truth. Its superior stability is highlighted in the corresponding MSE plot (Figure 11(d)), where the error stays bounded below an MSE of 0.004 for the majority of the horizon (up to 250 steps), peaking at just 0.006—less than half the maximum error of Model A. This demonstrates that Model B’s training strategy, Adams-Euler with AW3, more effectively mitigates the compounding error inherent in long-term auto-regressive forecasting.
@@
(c) Time-varying MSE of Model A
(d) Time-varying MSE of Model B
Figure 11: (a-b): $x$-velocity time series at seven probe points over 350 future snapshots. From bottom to top: probe points 1-7 (Figure 9). Solid lines: ground truth; circles: model predictions. (c-d): time-varying MSE for each model, averaged over the seven probe points.
@@ 4.5 Comparison with conventional noise-injection approach for long-term rollout
We evaluate three model configurations (direct prediction, forward Euler, and Adams-Euler time integration schemes) trained with Gaussian noise $\mathcal{N}(0,0.16^{2})$ injected into input data. This noise level, identified as optimal in our previous work [27], is consistently applied during training without multi-step rollouts.
The effectiveness of this approach is assessed using the temporal behavior of the $x$-velocity and the time-varying MSE at seven probe points downstream of the cylinder (Figure 12). The forward Euler model’s time-series plot (Figure 12(a)) shows significant deviations from the ground truth in both amplitude and phase, indicating that noise injection alone is insufficient for maintaining long-term accuracy. This instability is confirmed quantitatively by its time-varying MSE plot (Figure 12(c)), where the error grows rapidly to a maximum of approximately 0.1. In comparison, the Adams-Euler model, while still less accurate than our full framework, demonstrates greater stability in its time-series (Figure 12(b)). This is clearly evidenced by its corresponding MSE plot (Figure 12(d)), which shows the error peaking at just 0.0135—nearly an order of magnitude smaller than that of the forward Euler model. Comparing two schemes with noise injection, Adams-Euler (Figure 12(b)) still significantly outperforms forward Euler (Figure 12(a)), further demonstrating the robust performance of the Adams-Bashforth time integration scheme across various training approaches.
@@
(c) Time-varying MSE of forward Euler with noise injection model
(d) Time-varying MSE of Adams-Euler with noise injection model
Figure 12: (a-b) $x$-velocity time series at probe points over 350 future snapshots using noise injection. Solid lines: ground truth; circles: model predictions. (c-d): time-varying MSE for each model, averaged over the seven probe points.
Table 6 quantifies performance differences with MSE calculations for seven probe points. It’s important to note that noise injection itself provides substantial improvements over baseline approaches—comparing to Table 3 where no multi-step rollout or noise injection was applied (Direct: 0.125, Forward Euler: 0.138, Adams-Euler: 0.139), noise injection achieves significant performance gains (Direct: 0.019, Forward Euler: 0.017, Adams-Euler: 0.012), demonstrating its established effectiveness in the AR prediction community. However, for all time integration schemes, models trained with adaptive multi-step rollout consistently outperform even these improved noise injection results. The Adams-Euler scheme shows particularly dramatic enhancement, with the adaptive multi-step rollout (MSE of 0.002) achieving an 83% error reduction compared to the noise injection approach (MSE of 0.012)—representing not only a substantial improvement over baseline methods but also a significant advance beyond the already powerful noise injection technique.
Table 6: MSE comparison between adaptive multi-step rollout, noise injection, and a combined approach. Each multi-step rollout configuration uses the best adaptive weighting approach (footnoted). MSE values calculated from seven probe points in Figure 9.
@@
The performance gap stems from fundamental differences in addressing error accumulation. Noise injection builds resilience through input perturbations, making models more robust to small deviations, but does not explicitly address temporal dependencies and error propagation mechanisms inherent in AR predictions. In contrast, adaptive multi-step rollout enables learning from multiple future steps simultaneously, using strategically adjusted loss weights to focus on both immediate and distant predictions. This fundamental difference explains why adaptive multi-step rollout more accurately captures vortex shedding patterns compared to noise injection, as demonstrated in the time series comparisons (Figures 11 and 12).
We also investigated combining multi-step rollout with noise injection techniques, which revealed contrasting behaviors as shown in Table 6. The combined approach yields the best performance for direct prediction, while for derivative-based methods, this combination leads to training instability and significantly degrades performance. This contrasting behavior can be attributed to the interaction between the two regularization techniques. For direct prediction, the methods are complementary: noise injection robustifies the model to the exact type of input errors it will encounter during the multi-step rollout. Conversely, for derivative-based methods, the two techniques impose conflicting objectives. Multi-step rollout pushes the model to learn a precise, deterministic temporal trajectory. Noise injection, a form of stochastic regularization, forces the model to be robust to a distribution of inputs around that trajectory. When combined, noise is repeatedly propagated and potentially amplified through the unrolled computational graph of the multi-step loss. We hypothesize that this compounding of stochastic perturbations leads to unstable optimization process with higher MSE values.
4.6 Robustness evaluation under challenging conditions
To further validate our framework’s robustness, we evaluate its performance under three challenging conditions designed to simulate practical engineering constraints: (1) training with limited spatial information, where the model only sees a subset of the domain; (2) training with a larger time-step, which tests the numerical stability of the integration schemes; and (3) training on multiple flow scenarios simultaneously to assess the model’s generalization capability.
4.6.1 Performance under partial domain training
First, we evaluate performance under partial domain training, where models are trained on only a spatial subset ($0.3<x<0.75$ and $0.128<y<0.328$) of the original vortex shedding domain. This challenging scenario simulates practical applications with limited spatial coverage or memory constraints.
[…]
@@ 4.6.1 Performance under partial domain training
Quantitative analysis using seven probe points (Figure 14) confirms these results. Direct prediction completely fails to capture vortex shedding oscillations (MSE: 0.019), while forward Euler methods show gradual degradation (MSE: 0.011-0.013). Adams-Euler with AW3 achieves the lowest error (MSE: 0.008), accurately maintaining vortex shedding frequency, amplitude, and phase throughout the entire 350-step prediction horizon. These results validate the exceptional robustness of our Adams-Bashforth time integration with adaptive multi-step rollout framework, confirming its broad applicability for practical engineering applications where complete spatial information may be unavailable.
@@
4.6.2 Robustness to increased time-step size
A critical challenge for derivative-based prediction methods is their sensitivity to the temporal discretization, $\Delta t$. Because the model learns to approximate the temporal derivative based on a specific, fixed time step from the training data, the same $\Delta t$ must be used during the auto-regressive inference stage. To investigate this sensitivity and address concerns about the required temporal resolution, we conducted an additional ablation study to evaluate the framework’s stability under a coarser temporal resolution. For the cylinder flow case, where the original dataset has a fixed time step of $\Delta t=0.01s$, we retrained and evaluated the same four model configurations in Section 4.6.1 on a downsampled dataset with a doubled time step of $\Delta t=0.02s$. Since the total time period for training was kept constant, this effectively halved the number of snapshots available to the model.
The results, summarized in Table 7, reveal a stark degradation in performance for all models when trained on the coarser dataset. The MSE for all models increased by at least an order of magnitude. Our best-performing model (Adams-Euler + AW3) saw its MSE rise from 0.002 to 0.212, a hundred-fold increase. This failure is further evidenced by the highly inaccurate Strouhal number predictions. For instance, the direct prediction model predicted a frequency nearly double the ground truth ($St=0.2991$), while the derivative-based methods predicted frequencies less than half the true value ($St=0.065$).
Table 7: Time extrapolation performance comparison of top models when trained with the original time step ($\Delta t=0.01s$) versus a doubled time step ($\Delta t=0.02s$). Ground truth Strouhal number is 0.1438.
Model Configuration Original ($\Delta t=0.01s$) Doubled ($\Delta t=0.02s$)
MSE StSt MSE StSt
Direct + Vanilla Rollout 0.011 0.1407 0.0416 0.2991
Forward Euler + AW2 0.007 0.1489 0.1093 0.1516
Forward Euler + AW3 0.010 0.1407 0.2278 0.0656
Adams-Euler + AW3 0.002 0.1434 0.2120 0.0650
These findings underscore that while the choice of a time step can be critical, the key factor is whether the training data’s temporal resolution is sufficient to capture the core physical dynamics. In this case, halving the number of snapshots made the data too sparse for the models to learn the complex, periodic nature of vortex shedding. This suggests that the original $\Delta t=0.01s$ was already near the minimal sampling rate required. While it may be feasible to downsample a CFD dataset generated with an unnecessarily fine time step (e.g., one chosen for solver stability rather than to resolve physics), this experiment demonstrates that coarsening a dataset below the temporal resolution required to resolve its core physical dynamics will severely degrade prediction accuracy.
4.6.3 Generalization across multiple flow scenarios
To evaluate the framework’s ability to generalize, we again test the four leading model configurations explored in the previous robustness studies. For this experiment, each model was trained on a combined dataset comprising three distinct flow scenarios. Scenario 1 is the baseline case used throughout this paper, while Scenarios 2 and 3 are new additions with different inlet velocities, cylinder diameters, and mesh configurations (detailed in Figure 15 and Table 8). By exposing the models to more varied physical dynamics during training, this multi-scenario experiment tests their ability (AW3 especially, but including AW2 — both methods have tunable parameter $k$) to learn a more universal and robust representation of the underlying flow physics.
(a) Mesh for Scenario 2
(b) Mesh for Scenario 3
Figure 15: Two additional meshes used for evaluating the generalization performance across multiple flow scenarios.
Table 8: Physical and mesh properties of the three scenarios used for multi-scenario generalization testing.
Parameter Scenario Type
Scenario 1 Scenario 2 Scenario 3
Inlet $x$-velocity [m/s] 1.78 2.21 2.02
Cylinder diameter [m] 0.074 0.116 0.089
Number of nodes 1,946 1,852 1,925
Number of edges 11,208 10,644 11,082
Strouhal number 0.1438 0.1472 0.1490
The results in Table 9 highlight the robust performance of the Adams-Euler scheme combined with AW3 when trained on the aggregated multi-scenario dataset. While other models struggle to perform consistently across the varied dynamics, this configuration excels, achieving the lowest MSE for Scenarios 2 and 3 while maintaining strong performance on Scenario 1. Critically, its Strouhal number predictions (0.1462, 0.1441, 0.1428) are consistently accurate across all three scenarios, closely matching their respective ground truths (0.1438, 0.1472, 0.1490). This demonstrates its superior ability to learn from a diverse distribution of physical behaviors and accurately represent the distinct dynamics across different meshes, a key capability for developing more generalizable surrogate models.
Table 9: Time extrapolation performance evaluation (MSE and Strouhal number) of the four models trained on the combined multi-scenario dataset. The ground truth Strouhal numbers are 0.1438, 0.1472, and 0.1490 for Scenarios 1, 2, and 3, respectively.
Model Configuration Evaluated Scenario
Scenario 1 Scenario 2 Scenario 3
MSE StSt MSE StSt MSE StSt
Direct + Vanilla Rollout 0.035 0.0088 0.098 0.1462 0.076 0.1421
Forward Euler + AW2 0.028 0.1401 0.126 0.1387 0.078 0.1367
Forward Euler + AW3 0.024 0.1441 0.068 0.1428 0.053 0.1421
Adams-Euler + AW3 0.026 0.1462 0.051 0.1441 0.042 0.1428
@@ 5 Conclusions and Future Work
This study presents a comprehensive framework for enhancing long-term auto-regressive predictions in SciML models through the novel application of numerical time-integration schemes and adaptive multi-step rollout techniques. Our systematic evaluation across canonical 2D PDEs (advection, heat, and Burgers’ equations) first established a key hypothesis: as physical complexity increases, more sophisticated rollout techniques become essential for optimal performance. This trend was decisively validated in our most challenging test case of complex Navier-Stokes dynamics. For this system, our most advanced adaptive weighting strategies (AW2/AW3) proved crucial for achieving robust, long-term accuracy, confirming the insight gained from the simpler systems. By combining the two-step Adams-Bashforth scheme with these adaptive strategies, our lightweight GNN model—containing only 1,177 trainable parameters—demonstrated meaningful effectiveness under harsh constraints, achieving accurate predictions of complex Navier-Stokes dynamics across 350 future time steps and reducing the mean squared error from 0.125 to 0.002. Overall, our integrated methodology delivers an 89% improvement over fixed-weight multi-step rollout approach (reducing MSE from 0.018 to 0.002) and outperforms standard noise injection by 83% (reducing MSE from 0.012 to 0.002), while maintaining robustness even on truncated meshes. This powerful yet resource-efficient framework is designed to be model-agnostic, ensuring these advancements can benefit diverse scientific domains without specialized adaptations.
Several promising directions can emerge for future research. First, investigating higher-order time integration schemes beyond the two-step Adams-Bashforth could further enhance prediction stability, particularly for systems with complex temporal dynamics. Second, developing more computationally efficient multi-step rollout strategies would reduce the gradient computation overhead inherent in training across multiple future steps. Third, incorporating domain-specific physical principles—such as conservation laws from the Navier-Stokes equations—directly into the framework could enhance both accuracy and physical consistency while potentially reducing data requirements [35, 36]. While this study successfully demonstrates the framework’s effectiveness on 2D laminar flows, a crucial next step is to assess its scalability to more complex systems. Future work can therefore focus on extending the framework to three-dimensional simulations and turbulent flows, which may require architectural modifications and the integration of more advanced physics-informed constraints. Furthermore, exploring its applicability to multi-physics coupling problems represents another significant avenue for future research. Finally, a particularly promising direction is to integrate our framework with state-of-the-art architectures, including Fourier neural operators or Transolver [37], to validate its model-agnostic benefits and potentially push the boundaries of long-term prediction accuracy.